two buses leave Houston at the same time and travel in opposite directions. One bus averages 55 mph and the other bus averages 45mph. When will they be 400 mi apart?
Answer by oberobic(2304)

The distance formula is d=rt, where d is distance, r is rate (of speed0, and t = time.
d = 400
The buses are moving apart at the sum of their speeds:
r = 55+45
r = 100
.
Substituting what we know:
d = rt
400 = 100t
t = 4 hours
.
Checking the work:
Bus 1 will travel 55*4 = 220 miles in 4 hours
Bus 2 will travel 45*4 = 180 miles in 4 hours
220+180 = 400
So, in 4 hours they will be 400 miles apart.
.
Done.